# Autoformalizing and Simulating Game-Theoretic Scenarios using LLM-augmented Agents

### v1

@@ Abstract.
Game-theoretic simulations are a versatile tool for exploring interactions of both natural and artificial agents. However, modelling real-world scenarios and developing simulations often require substantial human expertise and effort. To streamline this process, we present a framework that enables the autoformalization of game-theoretic scenarios using agents augmented by large language models (LLMs). In this approach, LLM-augmented agents translate natural language scenario descriptions into executable logic programs that define the rules of each game, validating these programs for syntactic accuracy. A tournament simulation is then conducted, during which the agents test the functionality of the generated games by playing them. When a ground truth payoff matrix is available, an exact semantic validation can also be performed. The validated games can then be used in further simulations to assess the effectiveness of different strategies. We evaluate our approach on a diverse set of 55 natural language descriptions across five well-known $2\times 2$ simultaneous-move games, demonstrating 96% syntactic and 87% semantic correctness in the generated game rules. Additionally, we assess the LLM-augmented agents’ capability to autoformalize strategies for gameplay.
Key words and phrases:
autoformalization, game theory, large language models
1. Introduction
@@
Game-theoretic simulations are a versatile tool for exploring diverse phenomena, ranging from individual human interactions and organisational dynamics to biological processes, robotics, and interactions among artificial agents. However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort. Consider, for instance, the problem of autonomous vehicles coordination: designing a simulation model involves specifying the rules of interaction, the objectives of each agent, the information available to them, and the strategies they may employ, all while accounting for uncertainties and external factors such as traffic laws and pedestrian behaviour; that model subsequently has to be implemented. Recent advancements in Large Language Models (LLMs) enable a new approach to streamline these tasks.
@@
While LLMs have been evaluated as agents in game-theoretic contexts Akata et al. (2023); Fan et al. (2023); Guo (2023); Lorè and Heydari (2023), inherent limitations such as hallucinations Achiam et al. (2023) and logical and arithmetical errors Imani et al. (2023) restrict their utility as direct rational decision-makers. An alternative approach is to leverage LLMs’ capabilities in format translation, including the translation of natural language into formal representations. This process, known as autoformalization Wu et al. (2022), has shown success in creating formal representations for mathematics Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022) and logic Yang et al. (2023); Pan et al. (2023); Cosler et al. (2023); Chen et al. (2023); Feng et al. (2023). In this work, we apply autoformalization to automatically generate formal representations of game-theoretic scenarios.
This work builds on our previous research Mensfelt et al. (2024), where we introduced a basic framework for autoformalizing natural language scenarios that can be modelled using $2\times 2$ simultaneous-move games. In the earlier work, a formal solver was used to validate the syntactic correctness of the generated formal game representations, while semantic correctness was evaluated manually.
In this work, we integrate the previously introduced autoformalization module into a modular agent model and propose a framework for simulating tournaments with these agents. This framework serves two main purposes. First, by running the tournament, we can automatically verify both the functionality of the generated code and its semantic correctness. This verification can be performed approximately, by comparing the agent’s total payoff with the target payoff, or exactly—when the ground truth payoff matrix is available—by checking whether each combination of players’ moves produces the intended payoff. Second, the framework’s modular agent design supports the autoformalization of strategies, enabling the exploration of how different strategies–whether autoformalized or predefined–perform in autoformalized game scenarios. This approach facilitates the evaluation of strategy efficiency across diverse game-theoretic contexts. The main contributions of this work are as follows:
@@
  • An agent model with formal game and strategy representation: We introduce an agent model that incorporates a formal representation of both games and strategies, along with an autoformalization module, developed in the previous work.
@@
  • Three-level agent validation: We provide a mechanism for validating agents for syntactic correctness (through formal solvers), functionality (by employing the code in gameplay), and semantic correctness (by comparing target and achieved payoffs).
  • A simulation framework for autoformalized scenarios: A framework is introduced for simulating autoformalized game scenarios, enabling the investigation and comparison of different strategies – both autoformalized and predefined.
@@
  • Experimental evaluation: We evaluate the framework, demonstrating high syntactic and semantic correctness in the context of $2\times 2$ simultaneous-move games. We also test the capability of the agents to autoformalize 5 strategies.
The remainder of this paper is structured as follows. Sect. 2 introduces the preliminaries, including foundational concepts in game theory, large language models, and general game playing. In Sect. 3 we provide an overview of the proposed framework, followed by Sect. 4, which describes the methods used to evaluate the framework. Sect. 5 presents and analyses the results. Sect. 6 reviews related work in this area. Finally, Sect. 7 summarizes the conclusions and presents directions for future work.
2. Preliminaries
2.1. Game Theory
Game theory is a versatile and popular mathematical modelling framework that captures interaction scenarios pf varying complexity across various disciplines. Rasmusen Rasmusen (2006) composes games using four main components:
@@ 2.1. Game Theory
  • Actions: mappings between states and outcomes are actions carried out by the player Osborne and Rubinstein (1994);
@@
We formally define games in this context as the function of a set of $n$ players $N$, a non-empty set of actions $A_{i},\forall i\in N$, and for each player a utility function $u_{i}:A\rightarrow\mathbb{R}$ mapping a player i’s action to the corresponding payoff $\pi_{i}$.
@@
Table 1. A general normal form game describing the four possible outcomes $W,X,Y,Z$ where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$
In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity. In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt games are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$. This allows us to use the well-known terminology from Axelrod’s tournaments to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
@@
2.2. Large Language Models
The rapid development of natural language processing (NLP) through the advancements of transformer architectures Gillioz et al. (2020) and pre-trained models Qiu et al. (2020) led to the emergence of Large Language Models. State-of-the-art (SOTA) LLMs are evolving at a swift rate, largely due to the progressive increase in model parameters and training dataset size Zhao et al. (2023). While SOTA LLMs are improving with general use cases, there are approaches that further improve response accuracy and quality. Fine-tuning is the process of training LLMs on a specific custom dataset, increasing the model’s performance in the specific application context captured by the custom data Ziegler et al. (2020). Chain-of-thought prompting guides LLMs to break down their reasoning into simpler steps, often allowing models to deal with more complex requests or specific contexts more accurately Wei et al. (2023).
OpenAI’s GPT-4 Omni OpenAI (2024), employed in this work, is a multi-modal SOTA model, that supports textual, visual and audio data as well. GPT4-o has been shown to perform exceptionally well on reasoning benchmarks Wang and Zhao (2024).
2.3. General Game Playing
@@
General game playing Genesereth et al. (2005) focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention. The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games Love et al. (2006). GDL focused on information games only, so it was extended in GDL-II Thielscher (2010) to cover n-player games with incomplete information and games in extensive normal form Thielscher (2011). GDL-II is based on the standard syntax and semantics of logic programming and characterised by the special keywords shown in Table 2.
role(R) R is a player
init(F) F holds in the initial position
true(F) F holds in the current position
legal(R, M) R can do move M in the current position
does(R, M) player R does move M
next(F) F holds in the next position
terminal the current position is terminal
goal(R, N) R gets N points in the current position
Table 2. Subset of GDL-II (uppercase letters denote variables, lowercase letters denote predicates/function symbols)
@@
A challenge with GDL systems is that learning without human guidance demands complex reasoning. Players must infer the possible actions of others, effectively analysing hypothetical game scenarios before making decisions. Action formalisms like the classical Situation Calculus McCarthy and Hayes (1981) have been developed for precisely this purpose. Formal schemes and inference methods are readily available for Situation Calculus Giacomo et al. (2010); Lesperance et al. (2024), while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages. One such scheme Schiffel and Thielscher (2011) shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents Scherl and Levesque (2003). Our game solver, presented later in section 3.2, is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
3. LLM-augmented Autoformalizing Agents
3.1. Overview
Figure 1. The overview of our agent model. Green dotted lines denote internal communication and solid black lines external communication.
@@ 3.1. Overview
The framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents. An overview of the agent model is presented in Fig. 1. A logic programming component implemented in Prolog consists of two modules: a Game module containing game-independent and game-dependent rules, and a Strategy module describing which action to select against an opponent. Together, these modules form a solver (detailed in Section 3.2) that encapsulates the game logic and determines the agent’s next move. A Python wrapper maintains the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state. This modular design allows for easy substitution of game-dependent rules and strategies, either through predefined configurations or by autoformalizing natural language descriptions. Agents can also be saved to and reloaded from files. The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/autoformalizing-agents.
3.2. Solver
Our solver is based on logic programming, as described in Mensfelt et al. (2024), but repeated here for completeness. It comprises of a game-independent part representing the rules for any extensive-form game, a game-dependent part defining the rules of a particular game using the predicates from the game-independent part, and a set of auxiliary predicates completing the definition of that particular game. We follow standard Prolog conventions to interpret a game: variables are indicated by uppercase letters, while predicates and function symbols by lowercase. The symbol :- is interpreted as if, and the symbol \+ signifies not (negation by failure). An underscore ’_’ represents a variable whose value is ignored within a definition. The game state is then a situation, initially denoted by a constant (e.g. s0). The binary function do(M, S) denotes the situation resulting from the executing move M in situation S, consistent with the Situation Calculus.
3.2.1. Game-independent part
We define all legal transitions of an extended-form game from an initial situation S to a final situation F as follows:
game(F,F):-
final(F).
game(S,F):-
\+ final(S),
legal(M,S),
game(do(M,S),F).
The game terminates when the final game situation F is reached. Otherwise, in a non-final situation S, the game accepts a legal move M, and the game continues in the next do(M,S) situation, until the final situation F is reached. To determine what holds in each legal situation, we use Situation Calculus, represented as:
holds(F, S):-
initially(F, S).
holds(F, do(M, S)):-
effect(F, M, S).
holds(F, do(M, S)):-
holds(F, S),
\+ abnormal(F, M, S).
holds/2 here is similar to true/1 in GDL, but with an additional parameter of the situation in which the fluent is true. It states that a fluent F holds in the initial situation (init/1 in GDL), a new fluent F is initiated by the effects of a move M executed in a situation S (next/1 in GDL), and a fluent F persists after a move is made, provided it is not abnormal; abnormal fluents do not persist (implicit in GDL). We also use rules of the form:
finally(F, S):- Conditions.
to return derived fluents F describing the result of the game, when the Conditions hold in the final situation S.
3.2.2. Game-dependent part
To represent a specific game we need to define game-dependent predicates for the initial state initial/1, the legal moves legal/2, what holds in the initial game situation via initially/2, the effects of a move on a situation via effect/3, what stops persisting in a situation after the execution of a move via abnormal/3, the final situation final/1, and the result of the game via finally/2 definitions. To exemplify these definitions, we show how to describe a PD game to our solver. The initial situation s0 is defined as:
initial(s0).
What holds in this initial situation we specify it as:
initially(player(p1), s0).
initially(player(p2), s0).
initially(role(p1,row), s0).
initially(role(p2,col), s0).
initially(control(p1), s0).
initially(control(p2), s0).
These assertions define first the player names represented by unique identifiers (p1 and p2), their roles (p1 is the row player, while p2 is the column player), and then the fact that initially either of them can the game next (via the control/1 fluent,as in GDL). What holds in the initial situation changes as a result of move being made in the game. We represent moves as terms of the form move(P, M), where P is a player, and M is a move. As in PD it is possible for any player to defect (’D’) or cooperate (’C’), we express this as:
possible(move(P,’D’),S):-holds(player(P),S).
possible(move(P,’C’),S):-holds(player(P),S).
It is then legal for a player to make a possible move if they have the control to execute it in the current situation:
legal(move(P,M), S):-
possible(move(P,M), S),
holds(control(P), S).
As an effect of a legal move M being made by a player P, we record in the next situation that the player did that move (effect/3 is like next/1 in GDL):
effect(did(P, M), move(P, M), S).
Once a legal move is executed, the player loses control, which we specify in our framework as:
abnormal(control(P), move(P, M), S).
When a player loses control cannot play a move until control is passed back. When each player makes a move from the initial situation, the resulting final situation can be determined as:
final(do(M2, do(M1, S))):-
ground(M2),
ground(M1),
initial(S).
Assuming the payoff matrix following the structure presented in Table 1 defined as:
payoff(’D’, D’, 1, 1).
payoff(’C’, D’, 0, 5).
payoff(’D’, C’, 5, 0).
payoff(’C’, C’, 3, 3).
the outcome of the game holds information about the actual moves made by the players, and their payoffs.
finally(outcome(P1,M1,U1,P2,M2,U2), S):-
holds(role(P1, row), S),
holds(did(P1, M1), S),
holds(role(P2, col), S),
holds(did(P2, M2), S),
payoff(M1, M2, U1, U2).
We can then extract specific outcome information, e.g. the player’s utility, through a goal/2 fluent (as in GDL) e.g.:
finally(goal(P1, U1), S):-
finally(outcome(P1,_,U1,_,_,_), S).
finally(goal(P2, U2), S):-
finally(outcome(_,_,_,P2,_,U2), S).
This finalises the definition of PD, enabling us to use the game description above to reason about the game’s dynamics. For instance, if player p1 wanted to analyse the game and identify the actions required to achieve a utility of 5, then it asks the query:
?- game(s0,F), finally(goal(p1,5),F).
Applying the PD payoff matrix, our solver produces two solutions for F, separated by semicolons (;). The result false indicates ”no further answers”:
F = do(move(p2,’C’),do(move(p1,’D’),s0));
F = do(move(p1,’D’),do(move(p2,’C’),s0));
false.
In the first solution, p1 acts first and p2 second, while in the second solution, the order is reversed. This outcome is expected, as both players have initial control in our game definition, resulting in both action order combinations being shown.
3.2.3. Strategy
Strategies are submitted separately from the game-dependent component. In the current implementation, they may rely on specific predicates – default_move/2, opposite_move/2, possible/2, and finally/2 – which are required from the game-dependent solver. The select/4 predicate is used to determine the next move M of player P against opponent O in game situation S, as illustrated in the tit-for-tat strategy example below:
select(P, O, S, M):-
\+ holds(last_move(O, _LMo), S),
@@ 3.2.3. Strategy
Here, initially P selects the default move (typically to cooperate), otherwise it mirrors the opponent’s move in the last round. More sophisticated strategies can be expressed. For example, below we show how to express the best response of a player for a game as:
select(P, O, S, M):-
\+ holds(last_move(O, _), S),
holds(default_move(P, M),S).
select(P, O, S, M):-
holds(last_move(O, LMo), S),
findall(Ui-Mi,
(game(S, F),
finally(outcome(P,Mi,Ui,O,Mo,Uo),F),
Ui >= Uo),
Options),
sort(0, @>, Options, Ranked),
highest(Ranked, M).
This second strategy differs from the previous one on tit-for-tat in that its second clause illustrates how to use our representation of games to perform complex reasoning with the game rules. This is done by ranking all legal future outcomes in a given situation according to he utility these have for the player and the opponent, and select the move that gives the highest utility for the player. This strategy also demonstrates the complexity of the autoformalization task.
3.2.4. Constraint checking

### v2

# GAMA: Generative Agents for Multi-Agent Autoformalization

@@ Abstract
Multi-agent simulations facilitate the exploration of interactions among both natural and artificial agents. However, modelling real-world scenarios and developing simulations often requires substantial expertise and effort. To streamline this process, we present a framework that enables the autoformalization of interaction scenarios using agents augmented by large language models (LLMs) utilising game-theoretic formalisms. The agents translate natural language descriptions of interactions into executable logic programs that define the rules of each game, ensuring syntactic correctness through validation by a solver. A tournament simulation then tests the functionality of the generated game rules and strategies. After the tournament, if a ground truth payoff matrix is available, an exact semantic validation is performed. We evaluate our approach on a diverse set of 110 natural language descriptions exemplifying five $2\times 2$ simultaneous-move games, achieving 100% syntactic and 76.5% semantic correctness in the generated game rules for Claude 3.5 Sonnet, and 99.82% syntactic and 77% semantic correctness for GPT-4o. Additionally, we demonstrate high semantic correctness in autoformalizing gameplay strategies. Overall, the results highlight the potential of autoformalization to leverage LLMs in generating formal reasoning modules for decision-making agents.
1 Introduction
@@
Multi-agent simulations provide insights into diverse phenomena, ranging from individual human interactions and organizational dynamics to biological processes, robotics, and interactions among artificial agents. However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort. Consider, for instance, the challenge of coordinating autonomous vehicles: creating a simulation model involves specifying the rules of interaction, the goals of each agent, the information available to them, and the strategies they may employ, among other factors. Once developed, this model must then be implemented. The emergence of Large Language Models (LLMs) enables a novel method for streamlining these tasks through autoformalization. We demonstrate it using scenarios that can be modelled as $2\times 2$ simultaneous-move games.
@@
Although LLMs have been evaluated as agents in game-theoretic contexts Akata et al. (2023); Fan et al. (2023); Guo (2023); Lorè and Heydari (2023), their inherent limitations—such as hallucinations Achiam et al. (2023) and logical or arithmetical errors Imani et al. (2023)—constrain their effectiveness as rational decision—makers in simulations. Rather than relying on LLMs as autonomous agents, a more promising approach is to harness their strengths in format translation, particularly in converting natural language into formal representations. This process, known as autoformalization Wu et al. (2022), has shown promise in formalizing mathematics Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022) and logic Yang et al. (2023); Pan et al. (2023); Cosler et al. (2023); Chen et al. (2023); Feng et al. (2023). The resulting formal representations can be utilized in combination with subsymbolic models within a neuro-symbolic framework. This approach, as demonstrated by recent advances such as AlphaGeometry, has the potential to achieve human-like problem-solving abilities in challenging domains like geometry Trinh et al. (2024).
In this work, we introduce GAMA: Generative Agents for Multi-Agent Autoformalization, a framework that leverages autoformalization to generate formal representations of strategic interactions. These representations serve as the reasoning module of agents deployed in simulations. This research builds upon our prior work Mensfelt et al. (2024), where we presented a basic framework for autoformalizing natural language interaction scenarios. In this paper, we integrate the previously developed autoformalization module into a modular agent architecture within a simulation framework. The main contributions of this work are as follows:
  1. I.
@@
    An agent model with formal game and strategy representation: We propose an agent model that incorporates formal representations of both games and strategies, along with the autoformalization module from our earlier work. Game rules and strategies can either be autoformalized or provided using predefined rules.
  2. II.
@@
    Three-level agent validation: We introduce a validation mechanism for agents at three levels: (1) syntactic correctness, using formal solvers; (2) functionality, by testing the code in the simulation; and (3) semantic correctness, by comparing target and achieved payoffs (when a ground truth payoff matrix is available) or verifying compliance with game type constraints (when a ground truth game type is provided).
  3. III.
    A simulation framework for autoformalized scenarios: We present a simulation framework for autoformalized interaction scenarios, enabling automated validation and facilitating the exploration and comparison of different strategies for specific scenarios.
  4. IV.
@@
    Experimental evaluation: We evaluate the framework, demonstrating high levels of syntactic and semantic correctness for scenarios modelled as $2\times 2$ simultaneous-move games. Additionally, we assess the agents’ ability to autoformalize five distinct strategies.
2 Preliminaries
2.1 Game Theory
Game theory is an analytical framework for strategic interactions, where games are composed of four main components Rasmusen (2006):
@@ 2.1 Game Theory
  • Actions: mappings between states and outcomes, carried out players Osborne and Rubinstein (1994);
@@
We formally define games in this context as a tuple $(N,A,\Pi)$ consisting of a set of $n$ players $N$, a non-empty set of actions $A_{i},\forall i\in N$, and for each player $i$ a payoff function $\pi_{i}:\times_{j\in N}A_{j}\rightarrow\mathbb{R}$ mapping action profiles to player payoffs.
@@
Table 1: A general normal form game describing the four possible outcomes $W,X,Y,Z$, where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$.
In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity. In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$. This allows us to use the terminology from Axelrod’s tournaments Axelrod and Hamilton (1981) to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
@@
2.2 Large Language Models
The rapid advancement of natural language processing (NLP), driven by transformer architectures Gillioz et al. (2020) and pre-trained models Qiu et al. (2020), has led to the emergence of Large Language Models (LLMs). State-of-the-art (SOTA) LLMs continue to evolve rapidly, primarily due to the increasing scale of model parameters and training datasets Zhao et al. (2023).
In this work, we employ OpenAI’s GPT-4 Omni OpenAI (2024), a multimodal SOTA model capable of processing text, images, and audio. It has demonstrated strong performance on reasoning benchmarks Wang and Zhao (2024). Likewise, we utilize Anthropic’s Claude 3.5 Sonnet, another high-performing SOTA model, which excels across multiple benchmarks Anthropic (2024).
2.3 General Game Playing
@@
General game playing Genesereth et al. (2005) focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention. The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games Love et al. (2006). GDL focused on information games only, so it was extended in GDL-II Thielscher (2010) to cover n-player games with incomplete information and games in extensive normal form Thielscher (2011).
@@
A challenge with GDL systems is that learning without human guidance demands complex reasoning. Players must infer the possible actions of others, effectively analysing hypothetical game scenarios before making decisions. Action formalisms like the classical Situation Calculus McCarthy and Hayes (1981) have been developed for precisely this purpose. Formal schemes and inference methods are readily available for Situation Calculus Giacomo et al. (2010); Lesperance et al. (2024), while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages. One such scheme Schiffel and Thielscher (2011) shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents Scherl and Levesque (2003). Our game solver is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
3 Generative Agents for Multi-Agent Autoformalization
3.1 Overview
Figure 1: GAMA framework: each agent’s game rules and strategy can be predefined, autoformalized, or loaded from an existing agent. The initial agent pool undergoes syntactic validation, and agents that fail to produce syntactically valid code within a defined number of attempts are eliminated (denoted by a crossed circle). The remaining agents are validated in the tournament and subsequently undergo semantic validation.
@@ 3.1 Overview
The GAMA framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents (Fig. 1). An agent consists of an autoformalization module and a logic programming component implemented in Prolog, which is divided into two modules: a Game module, containing both game-independent and game-dependent rules, and a Strategy module, which describes the actions to select against an opponent. Together, these modules form a solver that encapsulates the game logic and determines the agent’s next move. A Python wrapper facilitates the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state. This modular design allows for easy substitution of game-dependent rules and strategies, either through predefined configurations or by autoformalizing natural language descriptions. The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/GAMA.
3.2 Solver
The solver comprises a game-independent part that represents the rules for any extensive-form game, a game-dependent part that defines the rules of a particular game using the predicates from the game-independent part, and a strategy part to select an action. Our solver is implemented in Prolog Körner et al. (2022); Wielemaker et al. (2012)—a declarative programming language well-suited for high-level problem definitions—but it can easily be substituted with another formal solver. The game-independent and game-dependent parts were described in detail in Mensfelt et al. (2024). To enable the agent incorporating the solver to play, we extended it with the strategy part. Additionally, we implemented semantic validation in Prolog using validity constraints. The strategy part and constraint-checking mechanism are described below.
3.2.1 Strategy
Strategies are submitted separately from the game-dependent component. The select/4 predicate is used to determine the next move M of player P against opponent O in game situation S, as illustrated in the tit-for-tat strategy example below:
select(P, O, S, M) if
not holds(last_move(O, _LMo), S) and
@@ 3.2.1 Strategy
Here, a player P selects the default move first (typically to cooperate), otherwise it mirrors the opponent’s move in the last round. holds/2 is based on the implementation of the Situation Calculus McCarthy and Hayes (1981). The strategy relies on specific predicates – default_move/2 and last_move/2 – which are required from the game-dependent part of the solver. More sophisticated strategies can be expressed.
3.2.2 Validity checking